how many litres of h20 is produced by the complete combustion of 2.00 mol of c4h10 assuming stp conditions
1
Expert's answer
2020-01-20T05:10:42-0500

2 C4H10 + 13 O2 -> 10 H2O + 8 CO2

2.00 mol C4H10 "\to" 10 mol H2O

m(H2O)=n*M= 10(mol)*18(g/mol)=180g

180g H2O=180mL H2O